OpenStudy (anonymous):

Find the sum of the convergent series \[\sum_{n=1}^{\infty} 1/n(n+3)\]

OpenStudy (anonymous):

Is it 11/6?

OpenStudy (anonymous):

11/18

OpenStudy (anonymous):

ok 11/18

OpenStudy (anonymous):

now let me check

OpenStudy (anonymous):

do you know how to write 1/n(n+3) in polynomial form?

OpenStudy (anonymous):

1/n^2+3n?

OpenStudy (anonymous):

no i mean it can be written as A/n - B/(n+3)

OpenStudy (anonymous):

oh partial fractions

OpenStudy (anonymous):

if you multiply the first part by n(+3) and the second by (n) you get A*(n+3)-B*n=1

OpenStudy (anonymous):

yeah partial fractions

OpenStudy (anonymous):

here A=1/3 and B=1/3

OpenStudy (anonymous):

B= - 1/3?

OpenStudy (anonymous):

forgot the negative on b

OpenStudy (anonymous):

And sum of ne from 1 to inf. (1/3 / n) - (1/3 / (n-3)

OpenStudy (anonymous):

I fyou start inserting numbers from 1; you get: 1/3[ 1+1/2+1/3+1+4+......) - 1/3( 1/4+1/5+1/6+.....)

OpenStudy (anonymous):

you see they cancel from 1/4 and only 1/3(1+1/2+1/3) remains

OpenStudy (anonymous):

and that is 11/18

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

telescope it until the terms can be cxl, then take the limit of the terms that are not numbers, then sum the numbers to get the final answer